Joseph mixed a certain amount of 10% isopropyl mixture with 6o gallons of a 20% isopropyl solution and the resulting mixture is 16% isopropyl. How much of the 10% isopropyl did he add?
Answer by rossiv53(27)

Ok, So let x = the amount of 10% mixture
10%x +20%(60) = 16%(x+60)
.10x +.20(60) = .16(x+60)
.10x + 12 = .16x+9.60
.10x-.16x+12 = .16x-.16x+9.60
-.06x+12 = 9.60
-.06x+12-12 = 9.60-12
-.06x = -2.40
x = 40
So, amount of 10% solution is 40 gallons
Good luck!